NottheMonga2087:

What is the solution to the linear equation? 6k + 10.5 = 3k + 12 k = 0.5 k = 2 k = 7.3 k = 9

supie:

\(6k+10.5=3k+12\) The first thing we do is subtract 3k from both sides which should give us \(3k=1.5\) because \(3k+10.5−10.5=12−10.5\) and 12-10.5=1.5 so we get \(3k=1.5\) then we ÷3 to both sides of the equation \(\LARGE\frac{3k}{3}=\frac{1.5}{3}\) \(\cancel{\frac{3k}{3}}=\frac{1.5}{3}\) (3k/3 cancels) all that is left is \(\frac{1.5}{3}\) so \(\frac{1.5}{3}=?\) k=?

NottheMonga2087:

k=5

supie:

Not quite-

supie:

\(1.5÷3=?\)

NottheMonga2087:

o wait

NottheMonga2087:

forgot to put the decimal

supie:

Yeah

NottheMonga2087:

0.5

supie:

Yes

NottheMonga2087:

k=0.5

NottheMonga2087:

thank you

supie:

Yes your answer is A

supie:

np